Two identical charges, 3.0 m apart, exert forces of magnitude 5.0 N on each other. What is the value of either charge?


1
Expert's answer
2022-03-02T14:41:11-0500

Force

"F=\frac{kq_1q_2}{r^2}"

"F=\frac{kq^2}{r^2}"

"q=\sqrt\frac{Fr^2}{k}"


"q_1=q_2=q=\sqrt\frac{5\times9}{9\times10^9}=70.71\mu C"